# LLMEval-3: A Large-Scale Longitudinal Study on Robust and Fair Evaluation of Large Language Models

### v3

@@ Abstract
Existing evaluation of Large Language Models (LLMs) on static benchmarks is vulnerable to data contamination and leaderboard overfitting, critical issues that obscure true model capabilities. To address this, we introduce LLMEval-3, a framework for dynamic evaluation of LLMs. LLMEval-3 is built on a proprietary bank of 220k graduate-level questions, from which it dynamically samples unseen test sets for each evaluation run. Its automated pipeline ensures integrity via contamination-resistant data curation, a novel anti-cheating architecture, and a calibrated LLM-as-a-judge process achieving 90% agreement with human experts, complemented by a relative ranking system for fair comparison. An 20-month longitudinal study of nearly 50 leading models reveals a performance ceiling on knowledge memorization and exposes data contamination vulnerabilities undetectable by static benchmarks. The framework demonstrates exceptional robustness in ranking stability and consistency, providing strong empirical validation for the dynamic evaluation paradigm. LLMEval-3 offers a robust and credible methodology for assessing the true capabilities of LLMs beyond leaderboard scores, promoting the development of more trustworthy evaluation standards. Our github repository is available at https://github.com/llmeval/llmeval-3.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has led to a proliferation of benchmarks designed to assess their capabilities (DBLP:journals/corr/abs-2307-03109; DBLP:conf/emnlp/LaskarABRKKJB0P24; DBLP:journals/corr/abs-2506-11094). However, these benchmarks predominantly rely on a static evaluation paradigm where models are tested on fixed, publicly available datasets (DBLP:journals/corr/abs-2503-04149). This approach is fundamentally vulnerable to data contamination and test set overfitting, contributing to a growing “evaluation crisis” where benchmark scores may no longer reliably reflect a model’s true, generalizable abilities (DBLP:journals/corr/abs-2412-03597; DBLP:conf/naacl/Deng0TGC24; DBLP:journals/corr/abs-2402-02823). For example, GPT-4 achieved exact match rates of 52% and 57% when guessing the masked options in MMLU (mmlu) test sets, far exceeding random chance (leakage:deng2024). Similarly, Qwen-1.8B has been shown to exactly replicate complete n-grams from both training and test splits of GSM8K (GSM8K) and MATH (MATH), including 25 exact 5-gram matches in the MATH test set, indicating potential undetected data leakage and memorization (benchmarkleakage:xu2024). Furthermore, static benchmark designs, dataset contamination, and biased evaluation protocols could create misleading perceptions of LLM capabilities, undermining the reliability of current performance assessments (DBLP:journals/corr/abs-2412-03597).
@@
The crisis compels us to shift our focus from what capabilities to evaluate, such as knowledge and reasoning, to a more foundational question: how to evaluate in a manner that is robust, fair, and resistant to strategic manipulation. Based on our analysis of the current evaluation crisis, we identify three fundamental challenges that must be addressed to construct a trustworthy evaluation framework. These challenges correspond to the core stages of any assessment: the data, the protocol, and the ranking system.
@@
Figure 1: The LLMEval-3 framework comprises three core stages. First, in Data Construction, diverse exam data are collected, filtered by experts, standardized into original questions in JSON format, and expanded to enhance formality and diversity. Second, Fair Testing involves sampling 1,000 unique questions for each model’s evaluation, transmitting data via a secure JWT stream, and collecting model responses. Third, Fair Ranking entails selecting evaluators using Cohen’s $\kappa$ coefficient under human-machine agreement, generating relative rankings based on evaluation scores (0–3), validating system robustness through ablation studies (incorporating varied reasoning styles, question resampling, and volume adjustments), and analyzing five key dimensions of model errors to verify ranking fairness. Column abbreviations: Eng. (Engineering), Econ. (Economics), Edu. (Education), Lit. (Literature), Mgmt. (Management), Sci. (Science), Hist. (History), Med. (Medicine), Mil. (Military).
@@
Challenge 1: How can we ensure the integrity of the evaluation data? The cornerstone of any benchmark is its test set. If the data is public or predictable, it becomes susceptible to leakage into model training sets, leading to inflated scores that reflect memorization rather than true capability. Therefore, building a benchmark that is inherently resilient to data contamination is the primary challenge. To address this, we construct a large-scale, private question bank of over 220k graduate-level questions, which are held privately and augmented with structural variations to prevent memorization.
@@
Challenge 2: How can we design an unpredictable assessment protocol? Even with private data, a static protocol with fixed test sets can be reverse-engineered or exploited over time. A truly robust evaluation requires a dynamic and unpredictable process that prevents any form of strategic gaming. To achieve this, we implement a dynamic evaluation protocol where models are served unseen, randomly sampled questions in each session through a secure, two-layer anti-cheating architecture, ensuring that each evaluation is unique and unpredictable.
@@
Challenge 3: How can we establish a fair and stable ranking system under dynamic conditions? When the evaluation content is not fixed, traditional absolute scoring becomes unreliable for cross-model comparison. A fair ranking system must remain stable and consistent, even when models are tested on different, albeit equivalent, sets of questions. To solve this, we develop a novel relative ranking system powered by a highly-calibrated LLM-as-a-Judge framework. This system establishes rankings by comparing model performance relative to each other within each evaluation session, rather than relying on absolute scores, ensuring fair and robust rankings that demonstrate negligible variance across different data samples.
@@
Collectively, these solutions constitute LLMEval-3, a comprehensive framework for dynamic LLM evaluation. We leverage this framework to conduct an extensive longitudinal study on our official platform, spanning from the second half of 2023 to the first half of 2025. Throughout this period, we have continuously evaluated nearly 50 leading proprietary and open-source models, typically assessing them within one week of their public release. To ensure the fairness and timeliness of our evaluation, we constantly update our private question bank. Each model undergoes at least three independent, randomly sampled evaluation runs to ensure the stability of our findings. To date, this rigorous, ongoing effort has accumulated over 150k evaluation data points.
@@
Leveraging this large-scale evaluation data, we conduct systematic investigations across three key research questions corresponding to our core challenges. Our empirical analysis reveals several key findings that provide new insights into LLM capabilities and evaluation practices. We summarize several key findings: we find that (a) all models converge to a performance ceiling around 90% with persistent gaps in specialized domains like literature and medicine, (b) dynamic rankings diverge significantly from static benchmarks, and static benchmarks suffer from severe data contamination, (c) our framework demonstrates exceptional ranking stability with negligible variance under multi-round resampling and varying sample sizes. More insightful findings are presented in Section 3.
@@
    We construct LLMEval-3, a large-scale anti-cheating evaluation platform featuring a proprietary 220k question bank, secure dynamic sampling protocols, and robust anti-manipulation mechanisms for trustworthy LLM assessment.
@@
    We conduct an extensive 20-month longitudinal evaluation campaign across nearly 50 leading models, accumulating over 150k evaluation data points through continuous anti-cheating assessments and comparative analysis with static benchmarks.
@@
    We conduct extensive empirical analysis across three research questions corresponding to our core challenges, revealing seven key findings about model performance ceilings, ranking stability, and contamination vulnerabilities in current evaluation practices.
@@ 2 Design
In this section, we detail the design and implementation of LLMEval-3, which addresses the three fundamental challenges through a three-stage framework. Dataset Construction tackles data integrity by building a contamination-resistant private question bank. Evaluation Process addresses unpredictable assessment through dynamic sampling and anti-cheating mechanisms. Ranking System implements a calibrated ranking system for fair model comparison.
@@ 2.1 Dataset Construction
As of early 2025, this process has resulted in the LLMEval-3 dataset, which comprises over 220k questions across six main categories. To maintain evaluation freshness and prevent contamination, we continuously expand the question bank through the same manual collection and automated augmentation pipeline. Detailed statistical analysis of the dataset, including disciplinary breakdown and content diversity, is provided in Appendix A.
@@ 2.2 Evaluation Process
Dynamic Question Sampling
To ensure unpredictability, each model evaluation is based on a unique set of 1,000 questions randomly sampled from our private question bank. A dynamic algorithm generates a unique, non-repeatable question sequence for each evaluation session, making it impossible to predict upcoming questions based on historical patterns. Furthermore, models are required to answer questions in the pre-allocated order, preventing any “cherry-picking” strategies. This dynamic sampling ensures that every evaluation is a distinct and isolated event, reflecting the model’s true generalization ability rather than its performance on a known set of problems.
Secure Anti-Cheating Architecture
@@ 2.3 Ranking System
LLM-as-a-Judge Evaluation
To quantify answer quality, we establish a standardized scoring metric with an integer range of [0,3], ensuring consistent evaluation of response efficacy across diverse model architectures. For scoring implementation, we uniformly employ GPT-4o (4o) as our judge, which has demonstrated high human-machine agreement through rigorous validation (detailed in Section 3.4). This choice of a single, validated scoring model eliminates potential biases introduced by varying evaluative criteria.
@@
Evaluation Metrics
To mitigate systematic bias introduced by random sampling questions, LLMEval-3 employs both relative score and absolute scores as evaluation metrics.
@@ 3 Experiment and Analysis
Based on the three core challenges identified in our introduction, we design the LLMEval-3 evaluation framework. In this section, we systematically investigate three critical research questions that arise from these challenges:
Research Question I: What authentic capability distributions and longitudinal trends do LLMs exhibit under LLMEval-3?
Research Question II: How does LLMEval-3 dynamic evaluation compare with static benchmarks regarding ranking accuracy and contamination issues?
Research Question III: How stable and reliable is LLMEval-3’s relative ranking system under multi-round resampling and human-machine consistency validation?
@@ 3.1 Experimental Setup
Benchmarking LLMs on LLMEval-3
We tracked over 50 LLMs from late 2023 to mid-2025, presenting full results in Appendix C and focusing here on 13 representative models (proprietary and open-source). Each model is evaluated across three prompting paradigms (Zero-Shot, Few-Shot, Chain-of-Thought) and 10 academic disciplines.
Case Study and Error Analysis
We sample incorrect responses from all evaluated models and manually classified failures into five categories: disciplinary knowledge, misunderstanding, logical reasoning, factual inaccuracies, and format compliance.
Ablation Studies
@@
Benchmark Comparison: We measure Spearman correlation between LLMEval-3 and static benchmarks (AGIEval (AGIEval), C-Eval (c-eval)) and perform fill-in-the-blank replay tests (1,000 questions, three attempts each) to assess contamination.
Ranking Validation: We conduct multi-round resampling (n=1000, 2000, 4000) to test stability. Human-machine agreement validation involves two independent rounds of human evaluation across 13 models, with Cohen’s $\kappa$ coefficients computed against three LLM-as-Judge evaluators. Second, we run an ablation study comparing our relative ranking to the traditional Elo scoring system.
@@ 3.2 Research Question I
Finding 1: All models converge to a performance ceiling of around 90% over a longitudinal period, with leading open-source LLMs rivaling proprietary SOTA.
Figure 2 illustrates the performance growth trajectories of different model series over time. Table 1 presents comprehensive evaluation scores and subject-level breakdowns for each model under the LLMEval-3 framework. We can find that nearly all models approach a performance ceiling around 90-93% on academic knowledge tasks. This convergence indicates fundamental limits in current model architectures for knowledge-intensive evaluation.
Model ZS FS CoT Mean Variance
Claude-4-Sonnet-Thinking 85.27 85.60 86.87 85.91 0.48
@@
Besides, the evaluation results show that top-performing models achieve remarkable scores: the proprietary Doubao-1.5-Thinking-Pro reaches 93.67 while the open-source DeepSeek-R1 achieves 91.23. These leading models substantially outperform established systems like GPT-4o at 82.51 and o3-mini at 78.80. This demonstrates that with sufficient scale and targeted fine-tuning, both proprietary and open-source LLMs can achieve comparable performance.
@@
Finding 2: Models demonstrate significant domain-specific performance variations, with specialized ’thinking’ abilities offering only marginal gains.
@@
Our analysis reveals a consistent pattern of domain-specific performance across all models. As shown in Table 1, models excel in technical and business domains like Management and Economics (achieving scores $\geq 9.40$) but consistently underperform in humanities and specialized fields such as Literature, Medicine, and Military (all scores below 9.0). This indicates robust mastery in some areas but also highlights persistent knowledge gaps. Further analysis into model variants shows that enabling specialized “thinking” modes provides only modest advantages. For example, Gemini-2.5-Pro-Thinking scores 91.00 overall (vs. 91.07 for the preview), while Claude-Sonnet-4-thinking outperforms its base version by just $\approx 0.03$ overall. These results suggest that while dedicated ’thinking’ optimizations offer slight cross-domain benefits, the primary performance differentiators remain rooted in the breadth of domain-specific knowledge, not enhanced reasoning modes.
@@
Finding 3: In dynamic knowledge-intensive evaluations, prompting paradigms have minimal impact, whereas external augmentation significantly boosts performance.
As shown in Table 2, the performance variance across Zero-Shot (ZS), Few-Shot (FS), and Chain-of-Thought (CoT) prompts remains below 1.6 points for all evaluated models, indicating that core capabilities are not significantly influenced by the prompting format. Detailed evaluation results for the other models are provided in Appendix C. In stark contrast, enabling web search boosts GPT-4o’s overall score by +1.23 (from 82.51 to 83.74), with the largest gains in knowledge-heavy domains like Medicine (+0.64) and Literature (+0.60). This demonstrates that while models possess a stable internal knowledge base, augmenting it with external information is a highly effective strategy for enhancing performance.
@@
Finding 4: Systematic error analysis reveals that disciplinary knowledge gaps and comprehension failures are the primary limitations of current models.
@@
Our comprehensive analysis of incorrect responses identifies disciplinary knowledge errors (47.7%) and misunderstanding errors (39.8%) as the two most prevalent failure modes, accounting for nearly 90% of all incorrect responses. This distribution suggests that current limitations stem primarily from knowledge gaps and comprehension deficiencies rather than systematic reasoning failures. For instance, Figure 3 illustrates characteristic failures where models either exhibit a clear lack of domain-specific knowledge or fail to capture critical contextual cues like the word “first”, instead matching only superficial details. These cases demonstrate that while models possess sophisticated language understanding, they still struggle with deep domain expertise and nuanced contextual interpretation, which defines their current performance boundaries.

### v5

# LLMEval-Fair: A Large-Scale Longitudinal Study on Robust and Fair Evaluation of Large Language Models

@@ Abstract
Existing evaluation of Large Language Models (LLMs) on static benchmarks is vulnerable to data contamination and leaderboard overfitting, critical issues that obscure true model capabilities. To address this, we introduce LLMEval-Fair, a framework for dynamic evaluation of LLMs. LLMEval-Fair is built on a proprietary bank of 220k graduate-level questions, from which it dynamically samples unseen test sets for each evaluation run. Its automated pipeline ensures integrity via contamination-resistant data curation, a novel anti-cheating architecture, and a calibrated LLM-as-a-judge process achieving 90% agreement with human experts, complemented by a relative ranking system for fair comparison. A 30-month longitudinal study of nearly 60 leading models reveals a performance ceiling on knowledge memorization and exposes data contamination vulnerabilities undetectable by static benchmarks. The framework demonstrates exceptional robustness in ranking stability and consistency, providing strong empirical validation for the dynamic evaluation paradigm. LLMEval-Fair offers a robust and credible methodology for assessing the true capabilities of LLMs beyond leaderboard scores, promoting the development of more trustworthy evaluation standards.
*Equal Contribution.
Corresponding Author.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has led to a proliferation of benchmarks designed to assess their capabilities (DBLP:journals/corr/abs-2307-03109; DBLP:conf/emnlp/LaskarABRKKJB0P24; DBLP:journals/corr/abs-2506-11094). However, these benchmarks predominantly rely on a static evaluation paradigm where models are tested on fixed, publicly available datasets (DBLP:journals/corr/abs-2503-04149). This approach is fundamentally vulnerable to data contamination and test set overfitting, contributing to a growing “evaluation crisis” where benchmark scores may no longer reliably reflect a model’s generalizable abilities (DBLP:journals/corr/abs-2412-03597; DBLP:conf/naacl/Deng0TGC24; DBLP:journals/corr/abs-2402-02823). For example, GPT-4 achieved exact match rates of 52% and 57% when guessing the masked options in MMLU (mmlu) test sets, far exceeding random chance (leakage:deng2024). Similarly, Qwen-1.8B has been shown to exactly replicate complete n-grams from both training and test splits of GSM8K (GSM8K) and MATH (MATH), including 25 exact 5-gram matches in the MATH test set, indicating potential undetected data leakage and memorization (benchmarkleakage:xu2024). Furthermore, static benchmark designs, dataset contamination, and biased evaluation protocols could create misleading perceptions of LLM capabilities, undermining the reliability of current performance assessments (DBLP:journals/corr/abs-2412-03597).
@@
The crisis compels us to shift our focus from what capabilities to evaluate, such as knowledge and reasoning, to a more foundational question: how to evaluate in a manner that is robust, fair, and resistant to strategic manipulation. Based on our analysis of the current evaluation crisis, we identify three fundamental challenges for constructing a trustworthy evaluation framework. These challenges correspond to the core stages of any assessment: the data, the protocol, and the ranking system.
@@
Figure 1: The LLMEval-Fair framework comprises three core stages. First, in Data Construction, diverse exam data are collected, filtered by experts, standardized into original questions in JSON format, and expanded to enhance formality and diversity. Second, Fair Testing involves sampling 1,000 unique questions for models’ evaluation, transmitting data via a secure JWT stream, and collecting model responses. Third, Fair Ranking begins by selecting evaluators based on human-machine agreement, measured by Cohen’s $\kappa$ coefficients. Subsequently, the process entails generating relative rankings, validating robustness through ablation studies, and analyzing model errors. Discipline abbreviations: Eng. (Engineering), Econ. (Economics), Edu. (Education), Lit. (Literature), Mgmt. (Management), Sci. (Science), Hist. (History), Med. (Medicine), Mil. (Military).
@@
Challenge 1: How can we ensure the integrity of the evaluation data? The cornerstone of any benchmark is its test set. If the data are public or predictable, they become susceptible to leakage into model training corpora, leading to inflated scores that reflect memorization rather than true capability. Therefore, building a benchmark that is inherently resilient to data contamination is the primary challenge. To address this, we construct a private question bank of over 220k graduate-level questions, augmented with structural variations to mitigate memorization.
@@
Challenge 2: How can we design an unpredictable assessment protocol? Even with private data, a static protocol with fixed test sets can be reverse-engineered or exploited over time. A truly robust evaluation requires a dynamic and unpredictable process that prevents strategic games. To achieve this, we implement a dynamic evaluation protocol where models are served unseen, randomly sampled questions in each session through a secure, two-layer anti-cheating architecture, ensuring that each evaluation is unique and unpredictable.
@@
Challenge 3: How can we establish a fair and stable ranking system under dynamic conditions? When the evaluation content is not fixed, traditional absolute scoring becomes unreliable for cross-model comparison. A fair ranking system must remain stable and consistent, even when models are tested on different, albeit equivalent, sets of questions. To solve this, we develop a novel relative ranking system powered by a highly-calibrated LLM-as-a-Judge framework. This system ranks models by comparing their performance within each evaluation session rather than relying on absolute scores, ensuring fair rankings with negligible variance across different data samples.
@@
Collectively, these solutions constitute LLMEval-Fair, a comprehensive framework for dynamic LLM evaluation. We leverage this framework to conduct an extensive longitudinal study on our official platform, spanning from the first half of 2023 to the second half of 2025. Throughout this period, we have continuously evaluated nearly 50 leading proprietary and open-source models, typically assessing them within one week of their public release. To ensure the fairness and timeliness of our evaluation, we constantly update our private question bank. Each model undergoes at least three independent, randomly sampled evaluation runs to ensure the stability of our findings. To date, this rigorous, ongoing effort has accumulated over 180k evaluation data points.
@@
Leveraging the large-scale evaluation data, we systematically investigate three research questions corresponding to our core challenges and find that (a) all models converge to a performance ceiling around 90% persistent gaps in specialized domains like literature and medicine, (b) dynamic rankings diverge from static benchmarks, and static benchmarks suffer from severe data contamination, (c) our framework demonstrates exceptional ranking stability with negligible variance under multi-round resampling and varying sample sizes. More insightful findings are presented in Section 3.
@@
    We construct LLMEval-Fair, a large-scale anti-cheating evaluation platform featuring a proprietary 220k question bank, secure dynamic sampling protocols, and robust anti-manipulation mechanisms for trustworthy LLM assessment.
@@
    We conduct an extensive 30-month longitudinal evaluation campaign across nearly 60 leading models, accumulating over 180k evaluation data points through continuous anti-cheating assessments and comparative analysis with static benchmarks.
@@
    We conduct extensive empirical analysis across three research questions corresponding to our core challenges, revealing eight key findings about model performance ceilings, ranking stability, and contamination vulnerabilities in current evaluation practices.
@@ 2 Design
In this section, we detail the design and implementation of LLMEval-Fair, which addresses the three fundamental challenges through a three-stage framework. Dataset Construction tackles data integrity by building a contamination-resistant private question bank. Evaluation Process addresses unpredictable assessment through dynamic sampling and anti-cheating mechanisms. Ranking System implements a calibrated ranking system for fair model comparison.
@@ 2.1 Dataset Construction
As of early 2025, this process has resulted in the LLMEval-Fair dataset, which comprises over 220k questions across six main categories. To maintain evaluation freshness and prevent contamination, we continuously expand the question bank through the same manual collection and automated augmentation pipeline. Detailed statistical analysis of the dataset, including disciplinary breakdown and content diversity, is provided in Appendix A.
@@ 2.2 Evaluation Process
2.2.1 Dynamic Question Sampling
To ensure unpredictability, each model evaluation is based on a unique set of 1,000 questions randomly sampled from our private question bank. A dynamic algorithm generates a unique, non-repeatable question sequence for each evaluation session, making it impossible to predict upcoming questions based on historical patterns. Furthermore, models are required to answer questions in the pre-allocated order, preventing any “cherry-picking” strategies. This dynamic sampling ensures that every evaluation is a distinct and isolated event, reflecting the model’s true generalization ability rather than its performance on a fixed test set.
2.2.2 Secure Anti-Cheating Architecture
@@ 2.3 Ranking System
2.3.1 LLM-as-a-Judge Evaluation
To quantify answer quality, we establish a standardized scoring metric with an integer range of [0,3], ensuring consistent evaluation of response efficacy across diverse model architectures. For scoring implementation, we uniformly employ GPT-4o (openai2024gpt4o) as our judge, which has demonstrated high human-machine agreement through rigorous validation (detailed in Section 3.4). This choice of a single, validated scoring model eliminates potential biases introduced by varying evaluative criteria.
@@ 2.3.1 LLM-as-a-Judge Evaluation
2.3.2 Evaluation Metrics
To mitigate systematic bias introduced by random sampling questions, LLMEval-Fair employs both relative score and absolute scores as evaluation metrics.
@@ 3 Experiment and Analysis
Based on the three core challenges identified in our introduction, we design the LLMEval-Fair evaluation framework. In this section, we systematically investigate three critical research questions that arise from these challenges:
Research Question I: What authentic capability distributions and longitudinal trends do LLMs exhibit under LLMEval-Fair?
Research Question II: How does LLMEval-Fair dynamic evaluation compare with static benchmarks regarding ranking accuracy and contamination issues?
Research Question III: How stable and reliable is LLMEval-Fair’s relative ranking system under multi-round resampling and human-machine consistency validation?
@@ 3.1 Experimental Setup
3.1.0.1 Benchmarking LLMs on LLMEval-Fair
We tracked nearly 60 LLMs from June 2023 to December 2025, presenting full results in Appendix E and focusing here on 17 representative models (proprietary and open-source). Appendix C provides a detailed overview of these representative models. Each model is evaluated across three prompting paradigms (Zero-Shot, Few-Shot, Chain-of-Thought) and 10 academic disciplines. In addition, we sample incorrect responses from all evaluated models and manually classify failures into five categories: disciplinary knowledge, misunderstanding, logical reasoning, factual inaccuracies, and format compliance.
3.1.0.2 Ablation Studies
@@
Benchmark Comparison: We measure Spearman correlation between LLMEval-Fair and static benchmarks (AGIEval (AGIEval), C-Eval (c-eval)) and perform fill-in-the-blank replay tests (1,000 questions, three attempts each) to assess contamination.
Ranking Validation: We conduct multi-round resampling (n=1000, 2000, 4000) to test stability. Human-machine agreement validation involves two independent rounds of human evaluation with Cohen’s $\kappa$ coefficients computed against three LLM-as-Judge evaluators. Second, we run an ablation study comparing our relative ranking to the traditional Elo scoring system. The introduction to the Elo scoring system can be found in Appendix D.
@@ 3.2 Research Question I
3.2.0.1 Finding 1: All models converge to a performance ceiling of around 90% over a longitudinal period, with leading open-source LLMs rivaling proprietary SOTA.
Figure 2 illustrates the performance growth trajectories of different model series over time. Table 1 presents comprehensive evaluation scores and subject-level breakdowns for each model under the LLMEval-Fair framework. Nearly all models approach a performance ceiling around 90 on academic knowledge tasks. This convergence indicates fundamental limits in current model architectures for knowledge-intensive evaluation.
Besides, the evaluation results show that top-performing models achieve remarkable scores: the proprietary Doubao-1.5-Thinking-Pro reaches 93.67 while the open-source DeepSeek-R1 achieves 91.23. These leading models substantially outperform established systems like GPT-4o at 82.51 and o3-mini. This demonstrates that with sufficient scale and training, both proprietary and open-source LLMs can achieve comparable performance.
Model Paradigm Statistic
ZS FS CoT Avg. Var.
Claude-Sonnet-4-Thinking 85.27 85.60 86.87 85.91 0.48
@@
3.2.0.2 Finding 2: Models demonstrate significant domain-specific performance variations, with specialized “thinking” abilities offering only marginal gains.
@@
Our analysis reveals a consistent pattern of domain-specific performance across all models. As shown in Table 1, models excel in technical and business domains such as Management and Economics, but consistently underperform in humanities and specialized fields including Literature, Medicine, and Military. This pattern suggests strong mastery in certain areas while highlighting persistent knowledge gaps elsewhere. Additional comparisons across model variants indicate that dedicated “thinking” modes yield only modest improvements. For instance, Claude-Sonnet-4-Thinking exceeds its base counterpart by approximately 0.03 points overall. Taken together, these results imply that the dominant drivers of performance differences are still domain coverage and specialized knowledge rather than enhanced reasoning modes.
@@
3.2.0.3 Finding 3: In dynamic knowledge-intensive evaluations, prompting paradigms have minimal impact, whereas external augmentation may boosts performance.
As shown in Table 2, performance varies only slightly across Zero-Shot (ZS), Few-Shot (FS), and Chain-of-Thought (CoT) prompting, with differences staying within three points for all evaluated models. This indicates that model capability is largely stable with respect to prompt format. Detailed results for the remaining models are reported in Appendix E. By contrast, external augmentation produces improvement. When web search is enabled, GPT-4o’s overall score increases by 1.23 points, rising from 82.51 to 83.74, with the largest gains appearing in knowledge-intensive domains such as Medicine and Literature. These results suggest that internal capability is relatively insensitive to prompting choices, whereas access to external information may strengthen performance in knowledge-heavy settings.
@@
3.2.0.4 Finding 4: Systematic error analysis reveals that disciplinary knowledge gaps and comprehension failures are the primary limitations of current models.
@@
Our error analysis shows that most failures are driven by missing domain knowledge and comprehension breakdowns. Disciplinary knowledge errors account for 47.7% of incorrect responses, while misunderstanding errors account for 39.8%, together comprising nearly 90% of all errors. This pattern indicates that current limitations are dominated by knowledge coverage and contextual understanding rather than systematic reasoning flaws. For example, Figure 3 presents representative cases in which models either lack the required domain-specific knowledge or miss decisive contextual cues, such as the keyword “first”, and instead match only superficial details. Overall, these cases highlight that although models exhibit strong language fluency, they remain constrained by gaps in specialized expertise and failures of nuanced contextual interpretation.